Status Report

NOAA GOES SATOPS Morning Report 15 November 2000

By SpaceRef Editor
November 15, 2000
Filed under

SATOPS MORNING REPORT – November 15, 2000


NOAA-16 rev 766 / W  1732Z on November 14:  Wallops ran this
pass in “normal” mode and took commanding authority for the duration of
the pass.  The SOCC test director at Wallops then carried out the
preplanned OV ground receiver command testing sequence.  All spacecraft
systems nominal.

NOAA-16 rev 769 / F at 2244Z on November 14:  VTX2 transmitter
was powered off.  This action was taken in response to the inability
to lock onto APT data both at the SOCC and in the user community as well
as a currently unexplainable thermal profile.  NOAA and NASA engineers
will continue to review data and will be in contact with Lockheed Sunnyvale
on Wednesday, November 15.

NOAA-15 rev 13029 / F 1827Z on November 14:  Commanded the solar
array offset to a +53 degrees.  The array was repositioned to it’s
intended position.  Other systems were nominal.

NOAA-14 rev 30284 / F 1543Z on November 14:  L1B noisy from spacecraft. 
Received 77.9% of the expected data.  Data recovery not possible due
to recorder management.


No change in the status of the DMSP spacecraft.

Fairbanks 13-meter antenna “C” is undergoing maintenance troubleshooting
until further notice.  Intermittent problems supporting DMSP spacecraft
have been noted recently and an attempt is being made to isolate the problem. 
DMSP supports have been moved to antenna “B”.

DMSP operations were nominal over the past 24 hours.


No change in the status of the GOES spacecraft.

GOES-10 at 0337Z on November 14:  Imager frame break.  2 lines
lost due to loss of signal sync.

GOES operations were nominal over the past 24 hours.


Questions or comments to

SpaceRef staff editor.